# Length-Controlled Margin-Based Preference Optimization without Reference Model

### v1

@@ Appendix B Downstream Result Analysis
To demonstrate the effectiveness of our method, we first adhere to established evaluation protocols and report the results of downstream tasks on the Hugging Face Open Leaderboard V1 for all models, as shown in Table 4. Overall, our findings indicate that the impact of our method varies across different tasks.
Minimal degradation in knowledge and reasoning abilities.
Compared to the SFT model and other preference optimization methods, our approach largely maintains MMLU performance with only a slight decline. This suggests that our method is effective in preserving both knowledge and reasoning capabilities.
Enhancement of Scientific and Commonsense Reasoning.
For ARC and HellaSwag benchmarks, our method generally improves performance compared to the SFT model and demonstrates competitive effectiveness relative to other preference optimization methods. This improvement can be attributed to the preference optimization dataset we used, which contains prompts related to scientific reasoning and commonsense reasoning—domains that closely align with these tasks. Consequently, our method enhances the SFT model’s capabilities in these areas.
Enhancement of Truthfulness.
For truthfulqa task, we find that our method improves TruthfulQA performance compared to the SFT model and nearly all other preference optimization methods. This improvement can be attributed to the preference optimization dataset, which includes instances that emphasize truthfulness. As a result, the model gains a better understanding of context and generates more truthful responses.
Decline in Mathematical Performance.
For the GSM8K task, our method leads to a decline in performance compared to the SFT model and other preference optimization methods. Notably, different preference optimization methods exhibit varying levels of success on this benchmark. We hypothesize that the removal of the reference model in our approach may result in a loss of capability for solving complex arithmetic problems. Given the difficulty of the GSM8K benchmark, several methods have been proposed to address this challenge. For instance, Step-DPO (Lai et al., 2024) treats individual reasoning steps as units for preference optimization rather than evaluating answers holistically, thereby enhancing the long-chain reasoning ability of LLMs.
In general, our method demonstrates a balanced trade-off in downstream performance. It effectively maintains general knowledge and reasoning abilities while enhancing scientific and commonsense reasoning, as well as truthfulness. However, it comes at the cost of reduced mathematical performance. These results suggest that the choice of preference optimization dataset plays a crucial role in shaping model capabilities. A deeper and more systematic investigation is necessary to fully understand the broader implications of preference optimization.
Appendix C Implementation Details
@@
For LMPO, we maintained a consistent batch size of 128 across all four experimental settings. The learning rates were configured as follows: 3e-7 for Mistral-Base (7B), 5e-7 for Mistral-Instruct (7B), 6e-7 for Llama-3-Base (8B), and 1e-6 for Llama-3-Instruct (8B). All models were trained for a single epoch using a cosine learning rate schedule with a 10% warmup phase. Optimization was performed using Adam (Kingma, 2014). Furthermore, the maximum sequence length was set to 1024 for Mistral-Base (7B) and 2048 for all other configurations. We use 42 as training random seed.
Hyperparameter in LMPO.
Table 5 outlines the hyperparameters used for LMPO across four different settings. For the parameter $\beta$, we follow the configuration from SimPO. Among these parameters, $h$, which represents the home-court advantage, typically requires more careful tuning. For $\lambda$, we set it to 1.0 for Mistral-Base and 0.2 for the other settings. As mentioned in the main article, selecting the appropriate value for $\lambda$ is crucial for LMPO performance.
@@
The hyperparameters utilized for evaluation in this study align with those adopted in SimPO.https://github.com/princeton-nlp/SimPO/tree/main/eval We sincerely appreciate the SimPO team for their generous contributions and invaluable insights.
@@
All training experiments reported in this study were performed on a system equipped with four A100 GPUs, following the procedures outlined in the alignment-handbook repository.https://github.com/huggingface/alignment-handbook

### v2

@@ Appendix C Downstream Result Analysis
To demonstrate the effectiveness of our method, we first adhere to established evaluation protocols and report the results of downstream tasks on the Hugging Face Open Leaderboard V1 and V2 for all models, as shown in Table 4.
Overview of LMPO Performance
The Language Model Preference Optimization (LMPO) method demonstrates remarkable effectiveness across diverse evaluation benchmarks when compared to alternative preference optimization approaches. Through careful analysis of the provided data, we can observe that LMPO achieves consistently strong results across different model architectures and benchmark categories. This method exhibits particular strengths in knowledge preservation, complex reasoning tasks, and mathematical problem-solving while maintaining competitive performance in truthfulness and common sense reasoning benchmarks.
Model Architecture Interactions and Performance Patterns
LMPO shows varied performance patterns across different model architectures and variants. When applied to base models, LMPO demonstrates exceptional effectiveness, achieving high rankings on both Mistral-Base and Llama3-Base variants. This suggests that LMPO is particularly adept at optimizing models without prior instruction tuning. For the Mistral-Base variant, LMPO excels in knowledge-intensive and reasoning-heavy tasks, achieving top scores on multiple benchmarks. Similarly, with Llama3-Base, LMPO leads in several key benchmarks. This consistent performance across diverse benchmarks indicates strong generalizability within base model architectures.
When applied to instruction-tuned models, LMPO maintains robust performance but with some variations. On Mistral-Instruct, LMPO achieves a top ranking among all methods, with particularly strong results on reasoning benchmarks and truthfulness evaluation. However, its performance on Llama3-Instruct is somewhat less consistent, ranking in the middle of the compared methods. While it still achieves best scores on several benchmarks, it demonstrates notably weaker performance on certain mathematical word problems compared to alternative approaches. This pattern suggests that LMPO’s effectiveness may vary depending on the underlying model architecture and prior tuning approach, with particular strengths in preserving core knowledge and reasoning abilities.
Performance Across Benchmark Categories
LMPO demonstrates distinctive performance patterns across different benchmark categories. In knowledge-intensive benchmarks such as MMLU-PRO, LMPO consistently achieves top performance across most model variants. This demonstrates LMPO’s strength in preserving and enhancing broad knowledge capabilities during preference optimization. For complex reasoning tasks represented by the BBH benchmark, LMPO shows consistently strong performance across all model variants, suggesting the method effectively optimizes for complex reasoning capabilities without compromising knowledge.
In mathematical reasoning tasks, LMPO displays an interesting dichotomy. It consistently performs exceptionally well on the MATH benchmark across all model variants, indicating particular effectiveness at preserving formal mathematical reasoning abilities. However, LMPO shows relatively weaker performance on the GSM8K mathematical word problem benchmark compared to other methods across all model variants. This suggests that while LMPO excels at formal mathematical reasoning, it may have specific limitations in optimizing for certain types of applied mathematical word problems or step-by-step reasoning tasks.
For language understanding and truthfulness benchmarks, LMPO demonstrates particularly strong performance on TruthfulQA for instruction-tuned models but more moderate results on base models. This suggests LMPO may be especially effective at enhancing truthfulness when applied to models with prior instruction tuning. On commonsense reasoning tasks like Winograd, LMPO shows variable performance across model variants, with stronger results on base models than instruction-tuned variants.
Comparative Analysis with Other Methods
When compared to other preference optimization approaches, LMPO demonstrates distinct strengths and limitations. Compared to KTO, LMPO generally performs better on knowledge-intensive tasks and formal mathematical reasoning, while KTO tends to achieve better results on mathematical word problems. Against SimPO, LMPO typically outperforms on knowledge tasks but shows weaker performance on instruction-following evaluations for most model variants. When compared to DPO, LMPO consistently shows stronger performance on knowledge benchmarks and mathematical reasoning, while DPO demonstrates competitive results on some instruction-following tasks but generally ranks lower overall.
The performance patterns across methods suggest that different preference optimization approaches may target different aspects of model behavior. LMPO appears to effectively preserve and enhance knowledge and reasoning capabilities while potentially having less impact on certain applied problem-solving skills, particularly in mathematical word problems. This indicates that the choice of preference optimization method should consider the specific downstream applications and tasks for which the model is intended.
Methodological Implications and Future Directions
The performance patterns observed for LMPO suggest several key methodological implications. First, LMPO effectively preserves model knowledge, as evidenced by strong performance on knowledge-intensive benchmarks. Second, it enhances reasoning capabilities, particularly in complex reasoning tasks and formal mathematical reasoning. Third, it improves truthfulness in instruction-tuned models, suggesting effective alignment with truthful responses. However, its consistent limitation in mathematical word problem solving represents a clear area for potential improvement.
LMPO represents a robust preference optimization method that performs particularly well on tasks requiring knowledge preservation and complex reasoning. Its effectiveness across different model architectures suggests it captures generalizable aspects of human preferences. Future work might focus on addressing the specific limitations in mathematical word problem solving while maintaining the method’s strengths in knowledge and reasoning tasks. Additionally, investigating the model-specific interactions could provide insights into how to further enhance LMPO’s effectiveness across different model starting points.
Appendix D Implementation Details
@@
For LMPO, we adopted a consistent batch size of 128 across all four experimental configurations. The learning rates were set as follows: 3e-7 for Mistral-Base (7B), 5e-7 for Mistral-Instruct (7B), 6e-7 for Llama-3-Base (8B), and 1e-6 for Llama-3-Instruct (8B). All models were trained for one epoch using a cosine learning rate schedule, incorporating a 10
Hyperparameters in LMPO.
Table 5 summarizes the hyperparameter settings used for LMPO across the four configurations. The value of $\beta$ follows the setup proposed in SimPO. Among the parameters, $h$ (representing the home-court advantage) typically requires more careful tuning. For the weighting factor $\lambda$, we set it to 1.0 for Mistral-Base and 0.2 for the other settings. As discussed in the main text, the choice of $\lambda$ plays a critical role in the effectiveness of LMPO.
@@
The evaluation hyperparameters used in this study are consistent with those adopted in SimPO.https://github.com/princeton-nlp/SimPO/tree/main/eval We are grateful to the SimPO team for their open-source contributions and valuable insights.
@@
All training experiments were conducted on a system equipped with four A100 GPUs. The experimental setup closely follows the procedures described in the alignment-handbook repository.§§§https://github.com/huggingface/alignment-handbook
Experimental issues.
Since our method builds upon SimPO and employs the same experimental setup, we primarily reference the results reported in SimPO. However, several researchers have noted in the GitHub issues of SimPO that they were unable to replicate the published results. To ensure a fair and rigorous comparison, we downloaded the official SimPO code and independently reproduced its experiments. Our results reveal that, for most models, the outcomes from the official implementation differ substantially from those presented in the original paper. Consequently, we report the results obtained through our independent reproduction.
Method DPO SimPO LMPO
Peak Memory (per GPU) 77 GB 69 GB 69 GB
Table 6: Peak GPU memory usage comparison. SimPO and DPO use 8×H100 GPUs; LMPO uses 4×A100 GPUs.
Appendix E Efficiency Analysis
The table 6 above summarizes the required RAM and provides a comparison among DPO, SimPO, and our proposed LMPO. It reports the peak GPU memory usage per device for SimPO and DPO in the Llama-3-Base setting with 8×H100 GPUs, while LMPO is evaluated using 4×A100 GPUs. Our LMPO implementation achieves approximately a 10% reduction in GPU memory consumption compared to DPO. Furthermore, although LMPO uses only half the number of GPUs employed by SimPO, it maintains an equivalent per-GPU memory footprint. These findings collectively demonstrate the superior memory efficiency of our approach.
For the Mistral-7B-Base model, following the default configuration, our method can run on devices equipped with four GPUs each having 48GB of memory (e.g., A40). This indicates that our approach has a relatively low dependency on RAM.